OpenStudy (anonymous):

What is the simplified form of 18g^6h^6 over 12g^9h^4?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

3 h^2 ---- 2 g^3

OpenStudy (anonymous):

@shaun786 and @joeywhite thanks. :))

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

could you help me with more?

OpenStudy (anonymous):

ok - one more - plz post others separately

OpenStudy (anonymous):

What is the simplified form of x^2 - 36 over x^2 - 12x + 36? (and okay) :)

OpenStudy (anonymous):

x^2 - 36 = (x + 6)(x - 6) and x^2 - 12x + 36 = (x - 6)(x - 6) can you continue?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Thanks so much by the way! :)

OpenStudy (anonymous):

the next step is (x + 6)(x - 6) ----------- (x - 6)(x - 6) notice ( x - 6) occurs inthe top and bottom of the fraction so can you tell me the next step in the simplifying process?

OpenStudy (anonymous):

Are you kidding? When it comes to math, I'm a total dope. I have tons of notes too. I just don't get it. I guess it's just not my strong spot.

OpenStudy (anonymous):

Thanks for your help though! :)

OpenStudy (anonymous):

the (x-6)'s cancel out: (x-6) divided by (x-6) is 1, as it is with anything else - anything divided by itself = 1 (except 0) so the final answer is (x +6) ----- (x - 6)

OpenStudy (anonymous):

ok - thast it